### v1

# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

Abstract
Retrieval-augmented generation (RAG) is widely used to ground Large Language Models (LLMs) for multi-hop question answering. Recent work mainly focused on improving answer accuracy via fine-tuning and structured or reinforcement-based optimization. However, reliable reasoning in response generation faces three challenges: 1) Reasoning Collapse. Reasoning in multi-hop QA is inherently complex due to multi-hop composition and is further destabilized by noisy retrieval. 2) Reasoning–answer inconsistency. Due to the intrinsic uncertainty of LLM generation and exposure to evidence–distractor mixtures, models may produce correct answers that are not faithfully supported by their intermediate reasoning or evidence. 3) Loss of format control. Traditional chain-of-thought generation often deviates from required structured output formats, leading to incomplete or malformed structured content. To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a Group Relative Policy Optimization (GRPO) based reinforcement learning framework that trains models to perform faithful reasoning during response generation. CRAFT employs dual reward mechanisms to optimize multi-hop reasoning: deterministic rewards ensure structural correctness while judge-based rewards verify semantic faithfulness. This optimization framework supports controllable trace variants that enable systematic analysis of how structure and scale affect reasoning performance and faithfulness. Experiments on three multi-hop QA benchmarks show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales, with the CRAFT model achieving competitive performance with closed-source LLMs across multiple reasoning trace settings.
1 Introduction
Retrieval-augmented generation (RAG) is widely used for multi-hop question answering (QA), where large language models (LLMs) generate answers grounded in retrieved candidate documents. A typical multi-hop QA task requires synthesizing information across multiple pieces of evidence that may be temporally, conceptually, or causally linked, while filtering out irrelevant or misleading distractors introduced by retrieval, calling for explicit reasoning traces that faithfully connect the supporting evidence to the final answer. As illustrated in Figure 1, reliable response generation in RAG systems faces three intertwined difficulties. First, multi-hop reasoning is inherently difficult, as it requires composing information across multiple dependent steps, and this difficulty is further amplified by noisy retrieval that introduces misleading or irrelevant documents. Second, generation by large language models is intrinsically uncertain, and under mixed evidence–distractor contexts, models may additionally exploit shortcut behaviors, producing answers that are weakly supported by the underlying reasoning or evidence. Third, models often fail to reliably follow required output formats during generation, leading to incomplete or malformed structured outputs and limiting controllability of the produced reasoning traces. These issues complicate the learning and assessment of robust reasoning in RAG systems.
Researchers have explored various approaches to improve accuracy in LLM-based multi-hop QA tasks, which can be broadly categorized along three complementary dimensions. (1) Explicit reasoning without guaranteed faithfulness. Early approaches encourage step-by-step reasoning through supervised fine-tuning (SFT), chain-of-thought prompting, and decomposition-based formulations Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Yao et al. (2022). While effective in improving answer accuracy, the resulting reasoning traces are not guaranteed to be grounded in the supporting evidence, especially under noisy retrieval, where models may bypass relevant documents and exploit spurious shortcuts. Trivedi et al. (2020); Press et al. (2023); Lanham et al. (2023); Turpin et al. (2023). (2) Post-hoc grounding without structured reasoning. Another line of work promotes evidence grounding by introducing citations, evidence constraints, or verification signals Gao et al. (2023); Wei et al. (2025); Li et al. (2025a); Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025). But these approaches often apply grounding after generation or only at the answer level, leaving intermediate reasoning traces free-form and unstructured. (3) Optimization of LLMs with coarse-grained rewards. More recent approaches explore reinforcement learning to optimize reasoning and answer generation in retrieval-augmented QA systems Asai et al. (2024); Li et al. (2025b); Chen et al. (2025). These methods typically rely on coarse outcome-level rewards tied to final answer correctness, providing limited supervision over intermediate reasoning and faithfulness properties, and allowing correct answers to be produced without aligned or faithful reasoning.
To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) for the response generation stage of RAG-based LLMs in multi-hop QA under noisy retrieved candidates. CRAFT trains generators to produce structured, machine-auditable reasoning traces, enabling systematic verification of reasoning faithfulness to retrieved evidence and generated answers. Our study consists of two aspects: 1) RL training with fine-grained reward supervision. Given a question and a retrieved candidate set containing both evidence and distractors, CRAFT optimizes reasoning trace generation using group-based reinforcement learning. Training is guided by a decomposed reward function that combines (i) deterministic rewards enforcing machine-checkable constraints, such as output structure validity, citation-set correctness when applicable, and answer matching, with (ii) judge rewards from a high-capacity LLM-as-a-judge that assess semantic faithfulness, including consistency across reasoning stages (e.g., planreasonanswer) and grounding to retrieved evidence. This design provides targeted supervision over unlabeled but critical semantic properties of multi-hop reasoning. 2) Controllable trace space and scale study. CRAFT defines a controllable space of reasoning trace variants by composing trace components (<plan>, <gold_docs>, <reason>, <answer>) along an auditability spectrum. This design enables systematic analysis of how trace structure influences performance and faithfulness, as well as how model scale interacts with trace complexity and audit constraints. Our contributions are three-fold:
Dual reward mechanisms for faithful multi-hop reasoning optimization. We propose CRAFT, a GRPO-based RL framework combining deterministic rewards for verifiable metrics and judge rewards for semantic faithfulness.
Systematic analysis of controllable reasoning traces. We study how different combinations of trace components affect model performance, and characterize how model scale shapes the preferred trace structures and their performance boundaries in small LLMs.
Experimental results demonstrate that CRAFT consistently improves answer accuracy and the faithfulness of generated reasoning traces across model scales. Notably, the CRAFT-trained 7B model achieves competitive performance with closed-source LLMs across different reasoning trace configurations on three multi-hop QA benchmarks.
2 Related Works
2.1 Faithful Reasoning in LLMs
Chain-of-thought (CoT) prompting Wei et al. (2022) improves reasoning in language models, but recent work questions whether these traces faithfully reflect the model’s actual decision process. Lanham et al. (2023) found that CoT faithfulness decreases with model scale, while Turpin et al. (2023) showed that CoT can be influenced by spurious input features; in safety-critical LLM-integrated systems, such unconstrained outputs undermine response reliability Zhang et al. (2025). Recent evaluation methods include Tutek et al. (2025)’s faithfulness measurement via unlearning reasoning steps, Arakelyan et al. (2025)’s FLARE framework using logic programming for verifiable reasoning, and Sui et al. (2025)’s approach anchoring responses to structured knowledge graph paths. However, these methods largely rely on post hoc analysis rather than enforcing machine-checkable structure during training.
2.2 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) Lewis et al. (2020) grounds language models in retrieved documents. For multi-hop reasoning, Trivedi et al. (2023) proposed IRCoT with interleaved retrieval, while recent work includes HopRAG Liu et al. (2025a) using passage graphs, DualRAG Cheng et al. (2025) combining reasoning-augmented querying with knowledge aggregation, and OPERA Liu et al. (2025b) coordinating planning and retrieval via RL. On generation, Gao et al. (2023) trained models for citations, while InstructRAG Wei et al. (2025) and RAG-DDR Li et al. (2025a) optimize generation with reward signals. Despite these advances, most approaches prioritize retrieval quality or answer-level grounding rather than producing structured reasoning traces.
2.3 Reinforcement Learning for LLMs
Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) aligns language models via policy optimization like PPO Schulman et al. (2017), DPO Rafailov et al. (2023), and RRHF Yuan et al. (2023). GRPO Shao et al. (2024) removes the need for a critic by normalizing rewards within sampled groups, enabling efficient training for reasoning models like DeepSeek-R1 Guo et al. (2025). For RAG, Chen et al. (2025) applied multi-agent RL, while R3-RAG Li et al. (2025b) jointly optimizes reasoning and retrieval with outcome and process rewards. However, most methods use scalar rewards, making it hard to provide targeted supervision over structured reasoning.
2.4 Multi-hop Question Answering
Multi-hop QA benchmarks include HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). Press et al. (2023) identified a compositionality gap in models, while Trivedi et al. (2020) showed they exploit shortcuts rather than genuine reasoning. Methods like Decomposed Prompting Khot et al. (2023), Least-to-Most Zhou et al. (2023), and ReAct Yao et al. (2022) improve performance. Recent work has focused on understanding reasoning capabilities: Yao et al. (2025) found that models can learn implicit multi-hop reasoning but require exponentially more training data as reasoning depth increases, while Yu et al. (2025) proposed back attention to enhance latent multi-hop reasoning. For evaluation, Lee et al. (2025) introduced GRADE, which generates difficulty-controlled QA pairs based on reasoning depth and retrieval difficulty. Despite these advances, most approaches lack machine-checkable grounding to verify that reasoning steps are supported by evidence.
3 Method
3.1 Problem Formulation
We study the response generation stage of RAG-based LLMs for multi-hop QA tasks, where each instance consists of a question and a retrieved candidate document set that mixes supporting evidence with distractors. Answering typically requires reasoning over multiple supporting documents indexed by , while distractors introduce spurious shortcuts. Our goal is to train models that generate not only correct answers but also auditable reasoning traces—structured outputs in XML-style format (<plan>, <gold_docs>, <reason>, <answer>) that expose the derivation process. Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding. This formulation transforms the learning problem from maximizing a sparse answer reward to optimizing a composite objective over the structure, process, and outcome of reasoning.
3.2 Overview
As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) a CRAFT Interface that defines structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
CRAFT Interface.
We define a family of XML-based output templates spanning an auditability spectrum. As detailed in Algorithm 1 (upper part), the full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning. Intermediate variants (CRAFT–CRAFT) selectively ablate <plan> or <gold_docs>, while CRAFT (Answer-Only) serves as a minimal baseline. This design enforces a chain of faithfulness: —any hallucination breaks at least one link and is detectable.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team (2024) | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
| Qwen2.5-1.5B Team (2024) | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
| Qwen2.5-3B Team (2024) | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
| Qwen2.5-7B Team (2024) | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
| Qwen3-4B Team (2025) | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
| Qwen3-30B Team (2025) | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
| API | ||||||||||
| GPT-5-mini OpenAI (2025) | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
| Deepseek-V3.2 DeepSeek-AI (2025) | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
| Gemini-2.5-Flash Google DeepMind (2025) | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team (2024) | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
| Qwen2.5-1.5B Team (2024) | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
| Qwen2.5-3B Team (2024) | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
| Qwen2.5-7B Team (2024) | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 7.25 | 13.13 | 0.80 | 24.80 | 34.61 | 5.40 | 22.85 | 28.46 | 1.85 |
| CRAFT | Trained | 19.35 | 27.99 | 28.75 | 45.70 | 60.01 | 57.25 | 44.70 | 52.05 | 47.85 |
| CRAFT | Trained | 30.96 | 39.69 | 45.85 | 54.53 | 69.06 | 79.50 | 56.50 | 64.75 | 76.10 |
| CRAFT | Trained | 54.35 | 65.03 | 80.35 | 66.61 | 82.09 | 95.00 | 78.72 | 84.57 | 94.80 |
3.3 CRAFT Training
We align models using GRPO with dual reward mechanisms to maximize a composite objective that calibrates structure, process, and outcome.
Principle
We optimize the policy to generate traces that are not only correct but structurally valid and reason-faithful, effectively supervising the “chain of faithfulness”. For each input , GRPO samples a group of traces and normalizes rewards within the group:
where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing . The key design choice is the total reward , which we decompose into four components:
where weights control the trade-off between strictness and accuracy.
3.3.1 (1) Deterministic Rewards
Rewards of this type can be computed directly from the trace structure and gold labels without requiring an LLM judge.
Format Compliance ().
We enforce strict adherence to the XML schema to ensure parsability.
where is the set of strings containing all required tags for variant in correct order. We set all other reward terms to when .
Citation Reward ().
We encourage document-level citations that cover the supporting documents while avoiding distractors. When <gold_docs> is present, we parse the predicted citation set and compute an index-level F1 against the gold supports :
Answer Correctness ().
We target final task accuracy with soft-match supervision.
where parses the <answer> content from the trace and is the standard token-level F1 score after normalization. This term ensures that process constraints do not degrade final performance, providing a continuous gradient toward the gold answer.
3.3.2 (2) Judge Rewards
Rewards of this type rely on an LLM judge to assess semantic consistency and evidence grounding, supervising the reasoning process rather than only the outcome.
Faithfulness Audit ().
We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1 for the detailed procedure). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
by averaging multiple machine-checkable criteria:
where and depends on the trace variant (e.g., plan/citation checks apply only when the corresponding fields exist). Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in .
Why no direct supervision for Plan/Reason text?
Unlike evidence selection (discrete indices) or final answers (short spans), the intermediate <plan> and <reason> are open-ended natural language generations with high semantic variance. Direct supervision (e.g., via SFT on teacher traces) often leads to brittle imitation or format collapse on out-of-distribution data. Instead of enforcing a rigid “gold plan,” our rewards focus on functional correctness via consistency checks (): a plan is valid if it is followed; a reason is valid if it is grounded. This allows the model to discover its own optimal reasoning paths within the auditable structure.
4 Experiment
| Model | Version | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Qwen2.5-7B | ||||||||||
| CRAFT | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 | |
| CRAFT | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 | |
| CRAFT | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 | |
| CRAFT | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 | |
| CRAFT | 31.55 | 41.97 | - | 53.40 | 67.92 | - | 50.60 | 59.29 | - | |
| Qwen2.5-7B | ||||||||||
| CRAFT | 32.35(-4.5) | 43.38(-4.4) | 50.85(-7.4) | 57.75(+0.8) | 72.04(+0.5) | 83.50(-0.2) | 62.65(+4.95) | 71.32(+4.6) | 78.95(+2.1) | |
| CRAFT | 29.85(-4.5) | 41.17(-3.5) | 56.10(-9.4) | 55.70(+0.9) | 70.25(+0.5) | 86.70(-1.4) | 59.25(+3.2) | 67.73(+2.1) | 78.30(+1.1) | |
| CRAFT | 35.60(-3.4) | 47.66(-1.9) | 61.35(-7.4) | 59.25(+1.5) | 73.68(+1.6) | 88.25(-2.4) | 65.40(+4.75) | 74.33(+4.5) | 85.80(+1.6) | |
| CRAFT | 31.80(-6.6) | 42.71(-6.6) | 66.60(-10.0) | 57.90(-1.4) | 72.43(-2.0) | 89.35(-4.4) | 63.15(+2.4) | 72.00(+1.8) | 82.25(-3.9) | |
| CRAFT | 27.00(-4.6) | 37.82(-4.2) | - | 53.30(-0.1) | 68.12(+0.2) | - | 56.15(+5.6) | 63.79(+4.5) | - | |
| CRAFT w/o judge reward | ||||||||||
| CRAFT | 51.37(+14.5) | 62.09(+14.3) | 71.20(+13.0) | 63.62(+6.7) | 79.40(+7.9) | 91.80(+8.1) | 75.91(+18.2) | 81.96(+15.2) | 89.25(+12.4) | |
| CRAFT | 55.93(+21.6) | 66.00(+21.4) | 69.45(+4.0) | 64.70(+9.9) | 80.01(+10.3) | 93.20(+5.2) | 77.45(+21.4) | 83.47(+17.8) | 84.80(+7.6) | |
| CRAFT | 52.75(+13.8) | 62.71(+13.2) | 72.15(+3.5) | 64.18(+6.4) | 79.21(+7.1) | 93.80(+3.1) | 74.98(+14.3) | 81.48(+11.7) | 88.00(+3.8) | |
| CRAFT | 55.82(+17.5) | 65.24(+15.9) | 76.85(+0.2) | 65.15(+5.9) | 80.97(+6.6) | 95.20(+1.5) | 78.09(+17.3) | 83.76(+13.6) | 88.80(+2.6) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
| CRAFT (full) | ||||||||||
| CRAFT | 54.35(+17.5) | 65.03(+17.2) | 80.35(+22.1) | 66.61(+9.7) | 82.09(+10.5) | 95.00(+11.4) | 78.72(+21.0) | 84.57(+17.9) | 94.80(+18.0) | |
| CRAFT | 58.35(+24.0) | 68.29(+23.7) | 78.35(+12.9) | 67.28(+12.5) | 82.10(+12.4) | 95.20(+7.2) | 79.70(+23.7) | 85.56(+19.9) | 89.90(+12.7) | |
| CRAFT | 54.65(+15.7) | 64.59(+15.1) | 81.40(+12.7) | 66.09(+8.3) | 81.04(+8.9) | 95.70(+5.1) | 76.88(+16.2) | 83.34(+13.5) | 93.20(+9.0) | |
| CRAFT | 56.91(+18.6) | 66.47(+17.2) | 86.70(+10.1) | 66.51(+7.2) | 81.91(+7.5) | 96.00(+2.3) | 79.39(+18.6) | 85.14(+14.9) | 94.10(+8.0) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
4.1 Experimental Setup
Datasets.
We evaluate CRAFT on three multi-hop QA benchmarks: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions Press et al. (2023); Trivedi et al. (2020). For training, we construct a mixed dataset by sampling from all three benchmarks, comprising 20,000 entries in total. For evaluation, we test on 2,000 samples from each benchmark.
Baselines.
We compare against two model categories: (1) Open-source pre-trained models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) evaluated via in-context prompting; (2) Closed-source API models: GPT OpenAI (2025), DeepSeek DeepSeek-AI (2025), and Gemini Google DeepMind (2025). For our method, we train models at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO Shao et al. (2024).
Metrics.
We report three evaluation metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes the token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (as detailed in Algorithm 1), providing a complementary measure beyond answer correctness.
Implementation Details.
All experiments are conducted on 4 NVIDIA H100 GPUs. We use the Qwen2.5 series as base models and train with GRPO using decomposed rewards. Training hyperparameters and configurations are detailed in Appendix B.
4.2 Main Results
Table 1 presents results evaluating CRAFT against open-source pre-trained models, closed-source API models, and SFT baselines across three multi-hop QA benchmarks. Additional experiments in Appendix C.
CRAFT improves accuracy across models at all scales.
For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%. The transformation is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model—a 16 improvement, showing how CRAFT’s decomposed rewards unlock latent reasoning capacity in small models.
CRAFT improves faithfulness across models at all scales.
For the 7B model, CRAFT achieves +22.0%/+8.9%/+16.1% EM improvements over SFT on MuSiQue/HotpotQA/2WikiMHQA. While SFT often degrades Faithfulness compared to base models (50.85% vs. 58.25% for 7B on MuSiQue, a drop of -7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT. This divergence illustrates that SFT learns to imitate surface patterns, whereas CRAFT’s judge-based faithfulness reward encourages internally consistent, evidence-grounded traces.
CRAFT matches or exceeds closed-source API models.
54.35% EM on MuSiQue (vs. GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs. 61.30%/80.09%), and 78.72% EM on 2WikiMHQA (vs. 70.55%, +8.2%). This demonstrates that GRPO can train a 7B open-source model to match or exceed proprietary systems.
4.3 Ablation Study
We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation. In CRAFT, the trace template determines deterministic reward availability (, , ), so variants ablate both trace components and their corresponding rewards. Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>, disabling all deterministic rewards and judge reward. We further ablate the judge reward by training with and without . Table 2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
Removing structure degrades performance.
For base models, answer-only CRAFT achieves the lowest EM (31.55%/53.40%/50.60% on MuSiQue/HotpotQA/2WikiMHQA), underperforming CRAFT (38.35%/59.30%/60.80%) by 5.9%–10.2%. While GRPO improves CRAFT substantially to 56.25%/63.25%/67.15%, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%). Notably, CRAFT cannot be evaluated for Faithfulness as it lacks reasoning traces.
Trace components reveal accuracy-faithfulness trade-offs.
For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without explicit training. After GRPO training, this gap narrows substantially (CRAFT: 80.35%/95.00%/94.80% vs. CRAFT: 86.70%/96.00%/94.10%), while CRAFTv1/CRAFTv2 improve auditability via explicit planning and citation.
CRAFT outperforms SFT across all configurations.
CRAFT achieves gains across all variants: +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models. In contrast, SFT shows frequent degradation (up to -6.6% EM on CRAFT) and systematic Faithfulness drops (-7.4% to -10.0% on MuSiQue). For CRAFT, GRPO improves MuSiQue EM by +17.5% (36.85%54.35%) while SFT decreases it by -4.5% (36.85%32.35%).
Judge reward is essential for faithfulness.
Without judge reward, Faithfulness drops -8.9% to -9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable. This validates our dual reward design: deterministic rewards ensure structural correctness and answer accuracy, while judge rewards are essential for semantic faithfulness.
4.4 Training and Traces Analysis
| Question: Which member of the research project conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005 was Kevin R. Henke most critical of? |
| Gold Answer: Russell Humphreys |
| plan |
| Subquestion1: Which research project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005? |
| Subquestion2: Who was Kevin R. Henke most critical of in [ANSWER_0]? |
| gold_docs |
| [2, 7] |
| reason |
| Subquestion1: From document [2], it is clear that the RATE project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005. |
| Subquestion2: Document [7] states that Kevin R. Henke has been critical of the RATE project’s results, specifically mentioning that he has argued that Russell Humphreys, a young-earth creationist involved in the project, has made errors in his research. |
| answer |
| Russell Humphreys ✓ |
4.4.1 Training Dynamics
Training performance scales with model size. Figure 3 shows total reward rising monotonically with scale: 0.5B (), 1.5B (), 3B (), and 7B (). Across all scales, reward curves converge smoothly without oscillation, indicating stable GRPO optimization under our multi-reward design. Notably, the largest relative gain appears at 1.5B, while marginal improvements diminish at 7B, suggesting saturation once most reward components approach their upper bounds.
Sub-rewards reveal fast format learning and a faithfulness threshold. Format Reward () converges within 50 training steps across all scales (0.5B: ; 7B remains near throughout), confirming that structural constraints are quickly internalized once the output template is learned. Answer Reward () scales with model capacity, reaching final scores of 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B, while Gold_doc Reward () improves steadily (0.050.27; 0.120.66; 0.190.55; 0.760.92), reflecting stronger evidence grounding in larger models.
Faithfulness exhibits a clear capacity threshold. Faithfulness Reward () remains near zero for 0.5B () and reaches only modest levels at 1.5B (0.10), whereas 3B and 7B show sharp improvements (, ). This transition indicates that structured consistency auditing is substantially more demanding than surface format compliance or answer correctness, and only emerges once the model has sufficient capacity to internalize and satisfy judge-based semantic constraints.
4.4.2 Output Traces Case Study
We analyze the quality of generated reasoning traces to understand how CRAFT structures multi-hop reasoning. Table 3 presents a representative example using the CRAFT, demonstrating CRAFT’s systematic approach with explicit planning, gold document citation, step-by-step reasoning, and final answer extraction. Additional case study in Appendix A. About traces format analysis in Appendix C.3.
5 Conclusion
We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA. CRAFT employs a dual reward design that combines deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness. Experiments on three benchmarks show that CRAFT consistently improves both accuracy and faithfulness over base models and SFT, with CRAFT matching or exceeding closed-source API models. Ablation results highlight the importance of structured traces and judge rewards for learning faithful reasoning, while training dynamics analysis indicates that faithfulness learning requires sufficient model capacity.
Appendix A Faithfulness Case Study
We present two representative cases from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs, demonstrating how CRAFT training improves both answer correctness and reasoning faithfulness. Each case includes the complete judge evaluation across all four faithfulness dimensions defined in Algorithm 1.
A.1 Case 1: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates how an unfaithful and incorrect response is transformed into a faithful and correct one.
Question: Who won the Indy car race in the largest city of the state where the Oh Yeah album performer was born?
Gold Answer: Mario Andretti
| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Where is the Oh Yeah album from?
Subq2: What is the largest city and capital of that state? Subq3: Who won the Indy car race in that city? |
Subq1: Where is the Oh Yeah performer from?
Subq2: What is the largest city and capital of the state where the Oh Yeah performer is from? Subq3: Who won the Indy car race in the largest city and capital of that state? |
| Gold Docs | [5, 12, 14] | [11, 12, 1] |
| Reason | [Answer1] Charles Mingus, the creator of the Oh Yeah album, was born in Nogales, Arizona (doc [5]). [Answer2] The largest city and capital of Arizona is Phoenix (doc [12]). [Answer3] The Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and the winner of the Indy car race there in 2016 was Ryan Hunter-Reay (doc [1]). | Doc [11] indicates that the Oh Yeah performer, Charles Mingus, is from Nogales, Arizona. Doc [12] states that the capital and largest city of Arizona is Phoenix. Doc [1] mentions that the Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and it is an Indy car race where Mario Andretti won. |
| Answer | Ryan Hunter-Reay | Mario Andretti ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 1 - Plan steps followed in order | 1 - Plan followed: performer→state→city→winner |
| Receipt-Reason () | 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs [11,12,1] in gold_docs |
| Reason-Answer () | 1 - Answer matches reasoning conclusion | 1 - “Mario Andretti” supported by reasoning |
| Evidence Grounded () | 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified in cited docs |
| Overall () | 0.0 - Failures: doc_outside_gold_docs, evidence_not_supported | 1.0 - All checks passed |
A.2 Case 2: Base (EM=1, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates that correct answers can arise from unfaithful reasoning, and CRAFT enforces faithfulness even when accuracy is preserved.
Question: What is the average yearly income for a person of the same nationality as the creator of MacGruder and Loud?
Gold Answer: $72,641
| Aspect | Base Model (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the nationality of [Answer1]? Subq3: What is the average yearly income for a person of [Answer2] nationality? |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the average yearly income for a person of the same nationality as the creator? |
| Gold Docs | [4, 5, 2] | [8, 4] |
| Reason | [Answer1]: From document 4, we learn that the creator of MacGruder and Loud is Aaron Spelling. Document 5 tells us that Lorraine Crosby, known as “Mrs. Loud,” is English. Therefore, the nationality of the creator of MacGruder and Loud is English. Document 2 provides the average yearly income for Americans, which is the closest nationality we have data for. The mean household income in the United States was $72,641. | Doc [8] states that MacGruder and Loud was one of the few failures from Aaron Spelling’s production company. Doc [4] provides information about Aaron Spelling, indicating that he was an American film and television producer. Therefore, the creator of MacGruder and Loud is American. The average yearly income for an American, as mentioned in doc [2], is $72,641. |
| Answer | $72,641 ✓ (correct answer, but unfaithful reasoning) | $72,641 ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 0 - Identifies Aaron Spelling but concludes “English” | 1 - Plan followed: creator→nationality→income |
| Receipt-Reason () | 0 - Misuses doc [5] (Lorraine Crosby creator) | 1 - All cited docs [8,4] in gold_docs |
| Reason-Answer () | 0 - Derives “English” but uses US income | 1 - “$72,641” supported by reasoning |
| Evidence Grounded () | 0 - “Creator is English” unsupported | 1 - Spelling→American→US income verified |
| Overall () | 0.0 - Failures: plan_ignored, answer_not_supported, evidence_not_supported | 1.0 - All checks passed |
A.3 Summary
These cases illustrate two critical aspects of CRAFT’s faithfulness enforcement:
Case 1 shows that CRAFT corrects both unfaithful reasoning (citing documents outside gold_docs, hallucinating unsupported claims) and incorrect answers.
Case 2 demonstrates that answer correctness does not imply reasoning faithfulness. The Base model produces the correct answer through a flawed reasoning chain (confusing Lorraine Crosby with Aaron Spelling), while CRAFT ensures the answer is derived through a consistent, evidence-grounded process.
The judge evaluation tables reveal that CRAFT training optimizes all four faithfulness dimensions simultaneously: plan-reason consistency (), receipt-reason consistency (), reason-answer consistency (), and evidence grounding (). This multi-dimensional supervision prevents shortcut learning and ensures that models produce not just correct answers, but trustworthy reasoning traces.
Appendix B Implementation Details
Training Framework.
We implement CRAFT using MS-Swift Zhao et al. [2025], an efficient and scalable fine-tuning framework for large language models. MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning. All experiments are conducted on NVIDIA H100 96GB x4 GPUs.
Model Configuration.
We train four model scales based on the Qwen2.5-Instruct series: 0.5B, 1.5B, 3B, and 7B parameters. All models use bfloat16 precision with Flash Attention 2 for memory-efficient training. Table 6 summarizes the key hyperparameters.
| Hyperparameter | Value |
| GRPO Configuration | |
| Batch Generations Accum. | 4 4 64 |
| Effective Batch Size | 1,024 |
| Max Completion Length | 256 tokens |
| Optimization | |
| Learning Rate | 5e-6 |
| LR Scheduler | Cosine |
| Warmup Ratio | 0.1 |
| Infrastructure | |
| Precision | bfloat16 |
| Attention | Flash Attention 2 |
| vLLM Max Length | 16,384 tokens |
Judge Model.
For the faithfulness reward and evaluation, we employ Qwen3-Max Team [2025] as the LLM judge. This model provides reliable semantic consistency assessment while maintaining computational efficiency. Because Qwen3-Max already functions as the judge for reward computation and evaluation, it is excluded from all performance comparison experiments to avoid role overlap.
Reward Function Configuration.
The reward functions vary across prompt templates based on their structural components:
CRAFT/CRAFT: , , (includes <gold_docs> field)
CRAFT/CRAFT: , (no explicit document citation)
CRAFT: , (direct answer generation)
All reward weights are set to 1.0, treating each component equally. The faithfulness reward is computed via LLM judge during evaluation but not used as a training signal in the ablation experiments (Table 2, “w/o judge reward”).
Training Data.
Each prompt template uses a dedicated training dataset of 20,000 samples, formatted according to the template’s structural requirements. The data distribution is: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA samples. Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns that indicate diverse linguistic characteristics across the three multi-hop QA datasets.
Appendix C Additional Experiments
This section presents comprehensive results for prompt template variants CRAFT–CRAFT, including training dynamics and evaluation results across all model scales.
C.1 Training Dynamics
Figures 5–8 illustrate the training dynamics for CRAFT–CRAFT templates. Consistent with CRAFT (main paper Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
Template-specific patterns: (1) CRAFT (Figure 5) shows similar dynamics to CRAFT, with all four reward components (, , , ) contributing to optimization. (2) CRAFT/CRAFT (Figures 6–7) lack but maintain stable training through format and answer rewards. (3) CRAFT (Figure 8) uses only and —notably, the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
Faithfulness reward verification.
To verify that the low training faithfulness reward (Figure 3) reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics (Algorithm 1) on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0). Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%. This indicates that while 69.7% of correct predictions exhibit fully faithful reasoning, 30.3% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness. This test-time value is approximately six times higher than the training-time average of 0.11 observed for this model (Figure 3), confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, resulting in lower average rewards during training, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
| Qwen3-4B Team [2025] | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
| Qwen3-30B Team [2025] | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
| Qwen2.5-1.5B Team [2024] | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
| Qwen2.5-3B Team [2024] | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
| Qwen2.5-7B Team [2024] | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 6.50 | 12.76 | 9.85 | 20.60 | 29.86 | 23.70 | 22.65 | 28.79 | 15.90 |
| CRAFT | Trained | 19.26 | 27.66 | 26.67 | 45.27 | 59.20 | 51.64 | 45.45 | 52.40 | 38.78 |
| CRAFT | Trained | 32.05 | 41.48 | 59.40 | 55.69 | 70.06 | 89.80 | 54.70 | 63.17 | 78.90 |
| CRAFT | Trained | 58.35 | 68.29 | 78.35 | 67.28 | 82.10 | 95.20 | 79.70 | 85.56 | 89.90 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
| Qwen3-4B Team [2025] | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
| Qwen3-30B Team [2025] | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
| Qwen2.5-1.5B Team [2024] | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
| Qwen2.5-3B Team [2024] | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
| Qwen2.5-7B Team [2024] | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 4.75 | 10.09 | 7.95 | 20.20 | 30.14 | 20.85 | 20.50 | 25.65 | 12.50 |
| CRAFT | Trained | 18.30 | 26.89 | 22.57 | 44.18 | 57.84 | 55.25 | 42.86 | 48.34 | 42.29 |
| CRAFT | Trained | 34.11 | 43.71 | 61.20 | 57.24 | 71.15 | 92.60 | 54.90 | 63.47 | 80.10 |
| CRAFT | Trained | 54.65 | 64.59 | 81.40 | 66.09 | 81.04 | 95.70 | 76.88 | 83.34 | 93.20 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
| Qwen3-4B Team [2025] | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
| Qwen3-30B Team [2025] | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
| Qwen2.5-1.5B Team [2024] | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
| Qwen2.5-3B Team [2024] | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
| Qwen2.5-7B Team [2024] | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 8.45 | 15.11 | 25.40 | 25.95 | 36.20 | 40.70 | 24.70 | 30.94 | 26.50 |
| CRAFT | Trained | 16.41 | 23.27 | 29.57 | 43.68 | 57.79 | 65.66 | 44.22 | 51.33 | 48.96 |
| CRAFT | Trained | 32.71 | 43.14 | 71.40 | 56.11 | 70.63 | 94.30 | 55.68 | 64.18 | 82.00 |
| CRAFT | Trained | 56.91 | 66.47 | 86.70 | 66.51 | 81.91 | 96.00 | 79.39 | 85.14 | 94.10 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | |||
| EM | F1 | EM | F1 | EM | F1 | ||
| Open-source | |||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
| Qwen3-4B Team [2025] | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
| Qwen3-30B Team [2025] | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
| API | |||||||
| GPT-5-mini OpenAI [2025] | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
| SFT | |||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
| Qwen2.5-1.5B Team [2024] | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
| Qwen2.5-3B Team [2024] | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
| Qwen2.5-7B Team [2024] | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
| Ours (RL) | |||||||
| CRAFT | Trained | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 | 0.05 |
| CRAFT | Trained | 22.95 | 31.91 | 45.70 | 59.04 | 38.40 | 46.19 |
| CRAFT | Trained | 31.80 | 42.07 | 53.15 | 67.10 | 49.30 | 57.19 |
| CRAFT | Trained | 56.25 | 65.38 | 63.25 | 78.00 | 67.15 | 73.77 |
| Model | Type | CRAFT | CRAFT | CRAFT | CRAFT | CRAFT | |||||
| Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | ||
| Qwen2.5-0.5B | |||||||||||
| Avg. | 0.0 | 98.3(+98.3) | 1.1 | 99.2(+98.1) | 0.0 | 94.3(+94.3) | 1.0 | 99.4(+98.4) | 9.0 | 0.1(-8.9) | |
| HotpotQA | 0.0 | 99.6(+99.6) | 0.1 | 99.8(+99.7) | 0.1 | 97.5(+97.4) | 0.8 | 99.8(+99.0) | 5.7 | 0.1(-5.6) | |
| 2WikiMQA | 0.0 | 99.5(+99.5) | 3.2 | 99.7(+96.5) | 0.0 | 92.5(+92.5) | 2.3 | 99.8(+97.5) | 16.5 | 0.1(-16.4) | |
| MuSiQue | 0.0 | 95.8(+95.8) | 0.0 | 98.2(+98.2) | 0.0 | 92.8(+92.8) | 0.0 | 98.5(+98.5) | 4.9 | 0.0(-4.9) | |
| Qwen2.5-1.5B | |||||||||||
| Avg. | 0.0 | 99.3(+99.3) | 2.1 | 95.1(+93.0) | 0.2 | 99.5(+99.3) | 1.2 | 81.9(+80.7) | 0.8 | 100.0(+99.2) | |
| HotpotQA | 0.0 | 99.5(+99.5) | 2.9 | 98.9(+96.0) | 0.1 | 99.2(+99.1) | 1.0 | 92.9(+91.9) | 0.7 | 100.0(+99.3) | |
| 2WikiMQA | 0.0 | 99.4(+99.4) | 0.8 | 99.2(+98.4) | 0.0 | 99.5(+99.5) | 1.4 | 91.2(+89.8) | 1.4 | 100.0(+98.6) | |
| MuSiQue | 0.1 | 99.1(+99.0) | 2.5 | 87.4(+84.9) | 0.6 | 99.7(+99.1) | 1.2 | 61.6(+60.4) | 0.2 | 99.9(+99.7) | |
| Qwen2.5-3B | |||||||||||
| Avg. | 13.2 | 96.9(+83.7) | 50.3 | 98.8(+48.5) | 6.4 | 98.0(+91.6) | 15.6 | 98.1(+82.5) | 41.6 | 100.0(+58.4) | |
| HotpotQA | 14.8 | 98.2(+83.4) | 55.2 | 99.2(+44.0) | 4.3 | 99.4(+95.1) | 27.6 | 99.2(+71.6) | 32.8 | 100.0(+67.2) | |
| 2WikiMQA | 7.7 | 99.0(+91.3) | 33.8 | 99.4(+65.6) | 6.3 | 99.1(+92.8) | 8.2 | 98.8(+90.6) | 23.9 | 100.0(+76.1) | |
| MuSiQue | 17.2 | 93.5(+76.3) | 61.8 | 97.6(+35.8) | 8.7 | 95.6(+86.9) | 10.8 | 96.3(+85.5) | 68.1 | 100.0(+31.9) | |
| Qwen2.5-7B | |||||||||||
| Avg. | 94.1 | 99.3(+5.2) | 99.2 | 100.0(+0.8) | 94.4 | 97.2(+2.8) | 61.6 | 99.8(+38.2) | 15.4 | 100.0(+84.6) | |
| HotpotQA | 96.1 | 100.0(+3.9) | 99.5 | 100.0(+0.5) | 95.7 | 99.1(+3.4) | 49.5 | 100.0(+50.5) | 11.2 | 100.0(+88.8) | |
| 2WikiMQA | 95.4 | 99.1(+3.7) | 99.3 | 100.0(+0.7) | 96.0 | 97.0(+1.0) | 72.9 | 100.0(+27.1) | 9.6 | 100.0(+90.4) | |
| MuSiQue | 90.9 | 98.9(+8.0) | 98.9 | 100.0(+1.1) | 91.6 | 95.5(+3.9) | 62.5 | 99.4(+36.9) | 25.4 | 100.0(+74.6) | |
C.2 Evaluation Results
CRAFT Results.
CRAFT retains the same four-field structure as CRAFT but with simplified instructions. CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base). Notably, CRAFT surpasses all API models on MuSiQue and 2WikiMHQA while achieving competitive Faithfulness scores (78.35%–95.20%).
CRAFT Results.
CRAFT removes explicit document citation (<gold_docs>), testing whether models can maintain faithfulness without explicit grounding signals. CRAFT achieves 54.65% EM on MuSiQue, 66.09% on HotpotQA, and 76.88% on 2WikiMHQA. Despite lacking supervision, the model maintains high Faithfulness (81.40%–95.70%), suggesting that plan-reason-answer consistency provides sufficient structure for faithful reasoning.
CRAFT Results.
CRAFT further simplifies the template while maintaining structured reasoning. CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA. EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance when combined with appropriate reward design.
CRAFT Results.
CRAFT represents the minimal template with direct answer generation, lacking structured reasoning traces ( and not applicable). CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA. A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training. This validates the importance of trace structure for model capacity constraints.
CRAFT Template Sensitivity Analysis.
Our experiments reveal a striking template sensitivity pattern unique to the smallest model, with different failure modes for SFT and GRPO training. SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%). In contrast, GRPO rescues templates CRAFT–CRAFT but still fails on CRAFT. Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
This dichotomy reveals two key insights: (1) SFT cold-start is fragile for 0.5B models—templates CRAFT/CRAFT with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse. (2) GRPO’s exploration mechanism overcomes SFT’s brittleness—reinforcement learning enables the model to discover valid reasoning paths even on complex structures (CRAFT/CRAFT), as long as some structural scaffolding exists. (3) Template CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure—direct answer generation lacks sufficient guidance for stable learning.
The optimal template for 0.5B deployment is CRAFT (<plan>+<reason>+<answer> without <gold_docs>), which provides the best balance: enough structure to guide reasoning without exceeding representational capacity, and compatibility with both SFT initialization and GRPO refinement.
C.3 Output Traces Format Analysis
Table 11 presents a comprehensive analysis of format compliance across all CRAFT variants and model scales.
GRPO dramatically improves format compliance across all scales. Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates (CRAFT–CRAFT), while even 3B models only reach 6.4%–50.3% on average. In stark contrast, GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases. Even 7B base models, despite already achieving 90.9%–99.5% on complex templates (CRAFT–CRAFT), benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%). This demonstrates that reward successfully teaches models to follow structured output requirements across all capacity levels.
Format complexity exhibits scale-dependent learnability. For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a puzzling anomaly on CRAFT, where GRPO models (0.1%) underperform base (9.0%) by -4.9% to -16.4% across datasets. This suggests the model defaults to more complex formats learned during multi-template training. In contrast, larger models master all formats: 1.5B achieves +80.7% to +99.3% gains, 3B reaches +58.4% to +91.6% improvements, and 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
Format Error Analysis. Table 12 presents representative error patterns across all CRAFT variants. Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
| Version | Base Model Errors (7B) | GRPO Model Errors (7B) |
| CRAFT |
Missing closing tags (60%):
<answer>1958 Award missing </answer> Truncation (30%): ...grandmother of Henry St John ( response cut mid-sentence |
Empty answer (rare):
<answer> no content, missing </answer> |
| CRAFT |
Missing closing tags (70%):
<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): <reason>...Michael Rubbo </answer> wrong closing tag None bare text |
No errors found ✓ |
| CRAFT |
Missing <answer> block (45%):
<reason>...answer is American. </reason> entire <answer>...</answer> missing Truncation (40%): ...cannot determine based on response cut off |
Missing/malformed tags (3%):
<answer>Infanta Maria missing </answer> <answer>The Woman in Fifth </ truncated closing tag |
| CRAFT |
Missing closing tags (85%):
<answer>Russell Humphreys missing </answer> |
Extra text after tags (0.2%):
<answer>No</answer> (Note: The reasoning...) parenthetical note not allowed |
| CRAFT |
Wrong format (85%):
A Bug’s Life no XML tags [Beaumont, Texas] bracket format [No] should be <answer>No</answer> |
No errors found ✓ |
C.4 Human Annotation Validation
| Aspect | Both | Both | LLM | LLM | Agr. | |
| Pass | Fail | Human | Human | |||
| Plan-Reason | 388 | 83 | 18 | 11 | 94.2% | 0.81 |
| Evidence Ground. | 380 | 80 | 23 | 17 | 92.0% | 0.75 |
| Answer Deriv. | 400 | 80 | 10 | 10 | 96.0% | 0.86 |
| Gold Doc Citation | 370 | 80 | 30 | 20 | 90.0% | 0.70 |
| Overall | 385 | 80 | 20 | 15 | 93.0% | 0.78 |
To validate the reliability of our LLM-based faithfulness metric, we conducted a human evaluation study following established protocols Adlakha et al. [2024]; Zheng et al. [2023]. Two expert annotators independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT with CRAFT template, using the same four-aspect criteria as our LLM judge (Algorithm 1). We measure inter-annotator reliability using Cohen’s Kappa Artstein [2017]:
where is the observed agreement and is the expected agreement by chance. Here denotes both-pass cases, denotes both-fail cases, (LLM pass, human fail) and (LLM fail, human pass) are disagreements, and is the total sample size. Table 13 presents the detailed per-aspect results. The results demonstrate consistently high agreement across all four aspects, with the overall faithfulness achieving 93.0% agreement and between automatic and human evaluation, indicating substantial agreement Zheng et al. [2023]; Bavaresco et al. [2024]. Answer Derivation shows the highest agreement (), as the correctness of final answers is relatively unambiguous. Gold Document Citation has lower agreement (), reflecting occasional ambiguity in determining whether cited documents are strictly necessary. Disagreements (20–50 cases per aspect) primarily occurred in edge cases involving implicit reasoning or borderline evidence usage. This strong alignment confirms that our LLM-based faithfulness metric serves as a reliable and scalable proxy for human judgment.