### v2

# CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering

Abstract
Retrieval-augmented large language models, when optimized with outcome-level rewards, can achieve strong answer accuracy on multi-hop questions. However, under noisy retrieval, models frequently suffer from right-answer-wrong-reason failures: they may exploit spurious shortcuts or produce reasoning traces weakly grounded in the supporting evidence. Furthermore, the lack of structured output control prevents reliable auditing of the underlying reasoning quality. To address this, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework for the response generation stage of retrieval-augmented multi-hop question answering. CRAFT trains models to produce structured reasoning traces with configurable levels of auditability (e.g., by selectively retaining planning, evidence citation, or reasoning steps). Training combines two complementary forms of supervision: deterministic rewards enforce verifiable constraints, including format compliance, answer correctness, and citation-set validity, while a judge-based reward audits semantic faithfulness by evaluating reasoning consistency and evidence grounding. Experiments show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales. Notably, semantic judge-based rewards improve answer accuracy rather than compromise it, enabling CRAFT to achieve performance competitive with strong closed-source models.
CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering
Yu Liu1, 2††thanks: Equal contribution., Wenxiao Zhang311footnotemark: 1, Diandian Guo1, 2, Cong Cao1††thanks: Corresponding authors., Fangfang Yuan1, Qiang Sun3, Yanbing Liu1, 2, Jin B. Hong3, Zhiyuan Ma422footnotemark: 2 1Institute of Information Engineering, CAS, 2University of Chinese Academy of Sciences, 3The University of Western Australia, 4Huazhong University of Science and Technology caocong@iie.ac.cn, mzyth@hust.edu.cn
1 Introduction
Retrieval-augmented generation (RAG) grounds large language models (LLMs) in external documents and is widely used for multi-hop question answering (QA), where answers require composing information across multiple passages Lewis et al. (2020). However, a correct answer does not imply faithful reasoning: models can exploit spurious shortcuts—citing irrelevant documents, confusing entities, or following logically inconsistent chains—yet still reach the gold answer (Figure 1, top). Such right-answer-wrong-reason failures are invisible to standard metrics that evaluate only the final output. The problem is especially acute in multi-hop QA, which demands compositional inference over dependent steps (e.g., resolving a bridge entity and linking its attributes to downstream facts) and is further destabilized by distractors that share surface overlap with the query Trivedi et al. (2020); Press et al. (2023).
Prior work addresses aspects of this problem from three directions, but each leaves significant gaps. (1) Explicit reasoning without guaranteed faithfulness. Chain-of-thought and decomposition-based methods encourage step-by-step reasoning Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Yao et al. (2022), but the resulting traces are not grounded in the retrieved evidence; under noisy retrieval, models may bypass relevant documents and produce plausible-looking but unfaithful rationales Lanham et al. (2023); Turpin et al. (2023). (2) Post-hoc grounding without structured reasoning. Citation and verification mechanisms introduce evidence signals such as document-level citations or consistency checks Gao et al. (2023); Wei et al. (2025); Li et al. (2025a); Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025), yet these are typically applied after generation or enforced only at the answer level, leaving intermediate reasoning steps unstructured and difficult to audit at scale. (3) Outcome-level optimization without process supervision. Reinforcement learning approaches optimize reasoning and generation Asai et al. (2024); Li et al. (2025b); Chen et al. (2025), but most rely on coarse outcome-level rewards tied solely to final answer correctness, providing weak credit assignment for multi-hop traces and no supervision over intermediate reasoning faithfulness.
To address these limitations, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) Shao et al. (2024) for the response generation stage of retrieval-augmented multi-hop QA. CRAFT defines a family of structured output templates that decompose generation into auditable stages (planning, evidence selection, step-by-step reasoning, and answer extraction), forming trace variants with configurable levels of auditability (Figure 1, bottom). Within each variant, these stages form a chain of faithfulness where each component constrains the next, so that unfaithful reasoning is more likely to violate at least one verifiable link. Training combines two complementary forms of supervision: deterministic rewards enforce format validity, citation-set correctness, and answer matching, while a judge-based reward from a high-capacity LLM audits semantic faithfulness by verifying cross-stage consistency (planreason, citationreason, reasonanswer) and grounding individual claims against the retrieved evidence. This extends process-level supervision beyond rule-based objectives to open-ended semantic properties (e.g., reasoning consistency and claim verifiability). The judge operates on binary decomposed criteria to limit bias propagation, and its reliability is validated against human annotations (§4.5). Together, these designs enable the model to learn not just what to answer but how to reason faithfully. Our contributions are as follows:
We propose CRAFT, an RL framework for retrieval-augmented QA that generates controllable trace variants (CRAFT–CRAFT), improving performance and reasoning auditability.
We introduce a dual-reward design that combines deterministic rewards for verifiable constraints with judge-based rewards for reasoning faithfulness, extending process-level supervision from rule-based objectives to open-ended semantic properties.
Experiments across multiple model scales show that CRAFT substantially improves reasoning faithfulness after training, and further reveal that semantic judge-based rewards can improve answer accuracy rather than compromise it. CRAFT achieves faithfulness and accuracy comparable to larger closed-source models.
2 Related Work
2.1 Faithful Reasoning and Grounding in RAG
CoT prompting Wei et al. (2022) improves reasoning but may not faithfully reflect the model’s decision process Lanham et al. (2023); Turpin et al. (2023); Zhang et al. (2025). Post-hoc faithfulness methods Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025) do not enforce structure during training. RAG Lewis et al. (2020) grounds models in documents, with multi-hop extensions including IRCoT Trivedi et al. (2023), HopRAG Liu et al. (2025a), and OPERA Liu et al. (2025b). Generation grounding via citations Gao et al. (2023) or reward signals Wei et al. (2025); Li et al. (2025a) primarily targets retrieval quality or answer-level grounding rather than structured reasoning with training-time faithfulness enforcement.
2.2 Reinforcement Learning for Reasoning
RLHF Ouyang et al. (2022) aligns models via PPO Schulman et al. (2017), DPO Rafailov et al. (2023), and RRHF Yuan et al. (2023); GRPO Shao et al. (2024) removes the critic via group-wise normalization for reasoning models Guo et al. (2025). Process rewards improve exploration efficiency Setlur et al. (2025) but require domain-specific verifier data. For RAG, multi-agent RL Chen et al. (2025) and joint reasoning-retrieval optimization Li et al. (2025b) have been explored, yet most methods use scalar answer-correctness rewards. LLM-as-a-Judge Zheng et al. (2023) offers scalable reward signals but introduces biases Chen et al. (2024); Gu et al. (2024); mitigations include binary criteria, rubric-based scoring, and human validation Adlakha et al. (2024); Bavaresco et al. (2024). CRAFT adopts these strategies with binary pass/fail judgments across decomposed dimensions, complemented by deterministic rewards.
2.3 Multi-hop Question Answering
Multi-hop QA benchmarks Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022) require composing information across multiple documents, yet models exhibit a compositionality gap Press et al. (2023) and exploit shortcuts Trivedi et al. (2020). Decomposition-based Khot et al. (2023); Zhou et al. (2023) and agentic Yao et al. (2022) methods improve performance, while recent work examines implicit reasoning Yao et al. (2025), back attention Yu et al. (2025), and difficulty-controlled evaluation Lee et al. (2025). Most approaches lack machine-checkable grounding to verify reasoning faithfulness.
3 Method
3.1 Problem Formulation
We study the response generation stage of RAG-based LLMs for multi-hop QA tasks, where each instance consists of a question and a retrieved candidate document set that mixes supporting evidence with distractors. Answering typically requires reasoning over multiple supporting documents indexed by , while distractors introduce spurious shortcuts. Our goal is to train models that generate not only correct answers but also auditable reasoning traces — structured outputs in XML-style format (<plan>, <gold_docs>, <reason>, <answer>) that expose the derivation process. Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding.
3.2 Overview
As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) CRAFT Trace Variants that define structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
Trace Templates.
We define a family of XML-based output templates spanning an auditability spectrum. The full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning. Intermediate variants selectively ablate components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains only <reason>+<answer>, and CRAFT (Answer-Only) serves as a minimal baseline. This design enforces a chain of faithfulness: — any hallucination breaks at least one link and is detectable. The trace template also determines which rewards are applicable: variants containing <gold_docs> activate , while variants with reasoning fields activate .
3.3 CRAFT Training
We align models using GRPO with dual reward mechanisms to maximize a composite objective that calibrates structure, process, and outcome.
Principle.
We optimize the policy to generate traces that are not only correct but structurally valid and reason-faithful, effectively supervising the “chain of faithfulness”. For each input , GRPO samples a group of traces and normalizes rewards within the group:
where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing:
The key design choice is the total reward , which we decompose into four components:
where weights control the trade-off between strictness and accuracy.
3.3.1 Deterministic Rewards
Rewards of this type can be computed directly from the trace structure and gold labels without requiring an LLM judge.
Format Compliance ().
We enforce strict adherence to the XML schema to ensure parsability.
where is the set of strings containing all required tags for variant in correct order. We set all other reward terms to when .
Citation Reward ().
We encourage document-level citations that cover the supporting documents while avoiding distractors. When <gold_docs> is present, we parse the predicted citation set and compute an index-level F1 against the gold supports :
Answer Correctness ().
We target final task accuracy with soft-match supervision.
where parses the <answer> content from the trace and is the standard token-level F1 score after normalization. This term ensures that process constraints do not degrade final performance, providing a continuous gradient toward the gold answer.
3.3.2 Judge Reward
Rewards of this type rely on an LLM judge to assess semantic consistency and evidence grounding, supervising the reasoning process rather than only the outcome.
Faithfulness Audit ().
We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
by averaging multiple machine-checkable criteria:
where and depends on the trace variant (e.g., plan/citation checks apply only when the corresponding fields exist). Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in . Each check yields a binary score; is their average. The binary, dimension-decomposed design limits the scope for judge error compared to holistic scoring. The judge model (Qwen3-Max) is distinct from the trained policy to avoid self-enhancement bias. Its reliability is validated against human annotations (§4.5), and cross-judge evaluation with an independent model confirms that results are not specific to the choice of judge (Appendix C).
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
| Qwen2.5-1.5B | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
| Qwen2.5-3B | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
| Qwen2.5-7B | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
| Qwen3-4B | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
| Qwen3-30B | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
[…]
| GPT-5-mini | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
| Deepseek-V3.2 | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
| Gemini-2.5-Flash | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
[…]
| Qwen2.5-0.5B | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
| Qwen2.5-1.5B | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
| Qwen2.5-3B | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
| Qwen2.5-7B | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
[…]
| CRAFT | Trained | 7.25 | 13.13 | 0.80 | 24.80 | 34.61 | 5.40 | 22.85 | 28.46 | 1.85 |
| CRAFT | Trained | 19.35 | 27.99 | 28.75 | 45.70 | 60.01 | 57.25 | 44.70 | 52.05 | 47.85 |
| CRAFT | Trained | 30.96 | 39.69 | 45.85 | 54.53 | 69.06 | 79.50 | 56.50 | 64.75 | 76.10 |
| CRAFT | Trained | 54.35 | 65.03 | 80.35 | 66.61 | 82.09 | 95.00 | 78.72 | 84.57 | 94.80 |
| Model | Version | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| CRAFT | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 | |
| CRAFT | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 | |
| CRAFT | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 | |
| CRAFT | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 | |
| CRAFT | 31.55 | 41.97 | - | 53.40 | 67.92 | - | 50.60 | 59.29 | - | |
[…]
| CRAFT | 32.35(-4.5) | 43.38(-4.4) | 50.85(-7.4) | 57.75(+0.8) | 72.04(+0.5) | 83.50(-0.2) | 62.65(+4.95) | 71.32(+4.6) | 78.95(+2.1) | |
| CRAFT | 29.85(-4.5) | 41.17(-3.5) | 56.10(-9.4) | 55.70(+0.9) | 70.25(+0.5) | 86.70(-1.4) | 59.25(+3.2) | 67.73(+2.1) | 78.30(+1.1) | |
| CRAFT | 35.60(-3.4) | 47.66(-1.9) | 61.35(-7.4) | 59.25(+1.5) | 73.68(+1.6) | 88.25(-2.4) | 65.40(+4.75) | 74.33(+4.5) | 85.80(+1.6) | |
| CRAFT | 31.80(-6.6) | 42.71(-6.6) | 66.60(-10.0) | 57.90(-1.4) | 72.43(-2.0) | 89.35(-4.4) | 63.15(+2.4) | 72.00(+1.8) | 82.25(-3.9) | |
| CRAFT | 27.00(-4.6) | 37.82(-4.2) | - | 53.30(-0.1) | 68.12(+0.2) | - | 56.15(+5.6) | 63.79(+4.5) | - | |
[…]
| CRAFT | 51.37(+14.5) | 62.09(+14.3) | 71.20(+13.0) | 63.62(+6.7) | 79.40(+7.9) | 91.80(+8.1) | 75.91(+18.2) | 81.96(+15.2) | 89.25(+12.4) | |
| CRAFT | 55.93(+21.6) | 66.00(+21.4) | 69.45(+4.0) | 64.70(+9.9) | 80.01(+10.3) | 93.20(+5.2) | 77.45(+21.4) | 83.47(+17.8) | 84.80(+7.6) | |
| CRAFT | 52.75(+13.8) | 62.71(+13.2) | 72.15(+3.5) | 64.18(+6.4) | 79.21(+7.1) | 93.80(+3.1) | 74.98(+14.3) | 81.48(+11.7) | 88.00(+3.8) | |
| CRAFT | 55.82(+17.5) | 65.24(+15.9) | 76.85(+0.2) | 65.15(+5.9) | 80.97(+6.6) | 95.20(+1.5) | 78.09(+17.3) | 83.76(+13.6) | 88.80(+2.6) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
[…]
| CRAFT | 54.35(+17.5) | 65.03(+17.2) | 80.35(+22.1) | 66.61(+9.7) | 82.09(+10.5) | 95.00(+11.4) | 78.72(+21.0) | 84.57(+17.9) | 94.80(+18.0) | |
| CRAFT | 58.35(+24.0) | 68.29(+23.7) | 78.35(+12.9) | 67.28(+12.5) | 82.10(+12.4) | 95.20(+7.2) | 79.70(+23.7) | 85.56(+19.9) | 89.90(+12.7) | |
| CRAFT | 54.65(+15.7) | 64.59(+15.1) | 81.40(+12.7) | 66.09(+8.3) | 81.04(+8.9) | 95.70(+5.1) | 76.88(+16.2) | 83.34(+13.5) | 93.20(+9.0) | |
| CRAFT | 56.91(+18.6) | 66.47(+17.2) | 86.70(+10.1) | 66.51(+7.2) | 81.91(+7.5) | 96.00(+2.3) | 79.39(+18.6) | 85.14(+14.9) | 94.10(+8.0) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
Why no direct supervision for Plan/Reason?
Unlike (discrete indices) and (short spans), and are open-ended generations with high semantic variance, making direct supervision via teacher traces brittle and prone to format collapse. Instead, we supervise functional correctness: is valid iff follows it (); is valid iff it is grounded (). This lets the model discover optimal reasoning paths within the auditable structure without imitating a fixed trace.
4 Experiments
4.1 Experimental Setup
Datasets.
We evaluate CRAFT on three multi-hop QA benchmarks: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions. For training, we construct a mixed dataset of 20,000 entries sampled from all three benchmarks. For evaluation, we test on 2,000 samples from each benchmark.
Baselines.
We compare against: (1) Open-source models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) via in-context prompting; (2) Closed-source API models: GPT-5-mini OpenAI (2025), DeepSeek-V3.2 DeepSeek-AI (2025), and Gemini-2.5-Flash Google DeepMind (2025). We train CRAFT at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO.
Metrics.
We report three metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (Algorithm 1).
Implementation Details.
4.2 Main Results
Table 3.3.2 presents results across three multi-hop QA benchmarks.
CRAFT improves accuracy across all model scales.
For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%. The improvement is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model, a 16 improvement, demonstrating that CRAFT’s decomposed rewards are effective even at very small scales.
CRAFT improves faithfulness across all model scales.
While SFT often degrades Faithfulness compared to base models (e.g., 50.85% vs. 58.25% for 7B on MuSiQue, a drop of 7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT.
CRAFT matches or exceeds larger closed-source API models.
CRAFT achieves 54.35% EM on MuSiQue (vs. GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs. 61.30%/80.09%), and 78.72% EM on 2WikiMHQA (vs. 70.55%, +8.2%). This demonstrates that GRPO with decomposed rewards can train a 7B open-source model to match or exceed proprietary systems.
4.3 Ablation Study
We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation. Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>. Table 3.3.2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
Removing structure degrades performance.
Answer-only CRAFT achieves the lowest base EM (31.55%/53.40%/50.60%), underperforming CRAFT by 5.9%–10.2%. While GRPO improves CRAFT substantially, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%).
Simpler traces achieve higher base faithfulness but lower auditability.
For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without training. After GRPO, this gap narrows substantially, while CRAFT/CRAFT improve auditability via explicit planning and citation.
SFT degrades faithfulness despite improving format compliance.
CRAFT achieves +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models, while SFT shows frequent degradation (up to 6.6% EM, 10.0% Faithfulness on MuSiQue).
Judge reward is essential for faithfulness.
Without judge reward, Faithfulness drops 8.9% to 9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable. This validates our dual reward design: deterministic rewards ensure structural correctness, while judge rewards are essential for semantic faithfulness. Notably, adding the judge reward also improves EM (e.g., +3.0% on CRAFT MuSiQue), confirming that faithful reasoning supervision enhances rather than compromises answer accuracy.
4.4 Training Dynamics
Figure 3 shows the training dynamics of GRPO across model scales using CRAFT(full format).
Training performance scales with model size.
Total reward increases monotonically with scale, and all curves converge smoothly, indicating stable optimization under our multi-reward design. Larger models achieve both higher starting points and higher ceilings, with the 7B model reaching a total reward of 3.41 compared to 1.47 for 0.5B.
Format compliance is learned quickly; faithfulness requires sufficient model capacity.
converges within 50 steps across all scales, indicating that structural constraints are the easiest to satisfy. scales with capacity, reaching 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B; notably, 1.5B and 3B achieve similar answer accuracy but diverge sharply on faithfulness. remains near zero for 0.5B and reaches only modest levels at 1.5B, whereas 3B and 7B show clear improvements (, ). This indicates that faithfulness learning is substantially more demanding than format compliance or answer accuracy, and only emerges with sufficient model capacity.
CRAFT achieves near-perfect format compliance across all scales.
Full format compliance results across all variants are reported in Appendix B.4.
| Aspect | Both | Both | LLM | LLM | Agr. | |
| Pass | Fail | Human | Human | |||
| Plan-Reason | 393 | 78 | 15 | 14 | 94.2% | 0.81 |
| Evidence Gr. | 375 | 83 | 25 | 17 | 91.6% | 0.75 |
| Answer Deriv. | 410 | 71 | 9 | 10 | 96.2% | 0.86 |
| Gold Doc Cit. | 370 | 80 | 33 | 17 | 90.0% | 0.70 |
| Overall (Avg.) | 387 | 78 | 20 | 15 | 93.0% | 0.78 |
4.5 Human Validation of Judge Reliability
To validate our LLM-based faithfulness metric, two annotators with NLP research experience independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT using the same four-aspect binary criteria as the LLM judge (Algorithm 1).
We select MuSiQue as it is the most challenging benchmark, requiring up to four reasoning hops. Results show 93.0% overall agreement with Cohen’s Artstein (2017), indicating substantial agreement (Table 3). Answer Derivation achieves the highest agreement (), while Gold Document Citation shows lower agreement (), reflecting inherent ambiguity in whether cited documents provide essential versus supplementary support. These results confirm that the LLM judge aligns well with human judgment in absolute terms. We further re-evaluate all configurations with an independent judge (Gemini-2.5-Flash) in Appendix C; model rankings are perfectly preserved (Spearman = 1.0). Together, these two evaluations address the circularity of using Qwen3-Max for both training and evaluation: human validation confirms that the judge’s criteria reflect genuine faithfulness rather than idiosyncratic preferences, while cross-judge consistency confirms that the reported improvements are not artifacts of a specific evaluator.
| Aspect | Base (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
[…]
Subq1: Who is the creator?
Subq2: What is the nationality? Subq3: What is the avg income? |
Subq1: Who is the creator?
Subq2: What is the avg income for same nationality? |
[…]
| 0 - Concludes “English” despite identifying Spelling | 1 - creatornationalityincome | |
| 0 - Misuses doc [5] (Crosby creator) | 1 - All cited docs in gold_docs | |
| 0 - Derives “English” but uses US income | 1 - $72,641 supported by reasoning | |
| 0 - “Creator is English” unsupported | 1 - All claims verified | |
| 0.0 | 1.0 | |
4.6 Case Study
Table 4 presents a case from MuSiQue where the base model produces the correct answer through unfaithful reasoning (EM=1, =0), while CRAFT achieves both correctness and faithfulness (EM=1, =1). The base model answers “$72,641” correctly but confuses Aaron Spelling with Lorraine Crosby from a distractor document, concluding the creator is “English” yet using US income data. This is a typical distractor-induced entity confusion: the model latches onto a surface-similar passage and constructs a plausible-sounding but internally inconsistent chain. Notably, all four judge checks fail (====0), demonstrating how the chain of faithfulness catches errors that outcome-level metrics miss entirely. In contrast, CRAFT correctly traces SpellingAmericanUS income, citing only relevant documents and passing all four faithfulness checks. An additional case study where both accuracy and faithfulness are corrected is provided in Appendix D.
5 Conclusion
We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA. By combining deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness, CRAFT consistently improves both accuracy and faithfulness across model scales, with CRAFT matching or exceeding closed-source API models. Notably, semantic judge-based rewards improve answer accuracy rather than compromise it.
Limitations
Our human validation of judge reliability is conducted on 500 samples from MuSiQue only. While we select MuSiQue as the most challenging benchmark and supplement with cross-judge evaluation (Appendix C), extending human annotation to all three datasets would further strengthen the reliability analysis; we leave this to future work due to the cost of expert annotation on multi-hop reasoning traces. Second, due to the substantial computational cost of GRPO training with online judge inference, our experiments are limited to models up to 7B parameters trained on 20K samples from three multi-hop QA benchmarks (HotpotQA, 2WikiMHQA, MuSiQue). A central finding of this work is that faithfulness supervision can improve rather than compromise answer accuracy. We focus on establishing this finding within the training distribution first, as confirming that the accuracy–faithfulness synergy holds under controlled conditions is a necessary prerequisite before testing its generalization. Scaling to larger models and evaluating whether this synergy transfers to out-of-distribution QA domains remains important future work.
Appendix A Implementation Details
This section expands on the experimental setup described in §4.
Training Framework.
We implement CRAFT using MS-Swift Zhao et al. (2025), an efficient fine-tuning framework for large language models. MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning. All experiments are conducted on NVIDIA H100 96GB 4 GPUs. Table A summarizes the key hyperparameters.
Model Configuration.
We train four model scales based on the Qwen2.5-Instruct series: 0.5B, 1.5B, 3B, and 7B parameters. All models use bfloat16 precision with Flash Attention 2 for memory-efficient training. For SFT baselines, we fine-tune each model on the same training data using standard cross-entropy loss with a learning rate of 2e-5 and 3 epochs.
| Hyperparameter | Value |
[…]
| Batch Generations Accum. | 4 4 64 |
| Effective Batch Size | 1,024 |
| Max Completion Length | 256 tokens |
[…]
| Learning Rate | 5e-6 |
| LR Scheduler | Cosine |
| Warmup Ratio | 0.1 |
[…]
| Precision | bfloat16 |
| Attention | Flash Attention 2 |
| vLLM Max Length | 16,384 tokens |
Judge Model.
We employ Qwen3-Max Team (2025) as the LLM judge for both the faithfulness reward during training and the Faithfulness metric during evaluation. The judge is prompted with the query, retrieved documents, and the model’s trace, and returns binary scores for each applicable audit dimension. We use temperature 0.0 for deterministic judge outputs. Because Qwen3-Max serves dual roles as reward provider and evaluator, it is excluded from all performance comparisons. We verify that this dual usage does not bias our conclusions through cross-judge evaluation with Gemini-2.5-Flash (Appendix C).
Reward Function Configuration.
The active reward components vary across trace templates based on the available structural fields: CRAFT/CRAFT: , , , (full supervision); CRAFT/CRAFT: , , (no citation reward as <gold_docs> is absent); CRAFT: , only (no reasoning fields for faithfulness audit). All reward weights are set to 1.0. This is a natural default since , , , and all produce values in ( is binary, and are F1 scores, and is an average of binary checks), so equal weighting does not introduce scale imbalance. Moreover, converges to 1.0 within 50 training steps across all scales (Figure 3), contributing negligible gradient signal thereafter. In the ablation without judge reward (Table 3.3.2, “w/o judge reward”), is removed from training but still computed during evaluation.
Training Data.
Each prompt template uses a dedicated training dataset of 20,000 samples: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA. We oversample MuSiQue as it contains questions requiring 2–4 reasoning hops, providing greater diversity in reasoning depth compared to the predominantly 2-hop questions in HotpotQA and 2WikiMHQA. Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns across the three datasets.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
| Qwen2.5-1.5B | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
| Qwen2.5-3B | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
| Qwen2.5-7B | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
| Qwen3-4B | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
| Qwen3-30B | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
[…]
| GPT-5-mini | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
| Deepseek-V3.2 | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
| Gemini-2.5-Flash | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
[…]
| Qwen2.5-0.5B | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
| Qwen2.5-1.5B | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
| Qwen2.5-3B | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
| Qwen2.5-7B | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
[…]
| CRAFT | Trained | 6.50 | 12.76 | 9.85 | 20.60 | 29.86 | 23.70 | 22.65 | 28.79 | 15.90 |
| CRAFT | Trained | 19.26 | 27.66 | 26.67 | 45.27 | 59.20 | 51.64 | 45.45 | 52.40 | 38.78 |
| CRAFT | Trained | 32.05 | 41.48 | 59.40 | 55.69 | 70.06 | 89.80 | 54.70 | 63.17 | 78.90 |
| CRAFT | Trained | 58.35 | 68.29 | 78.35 | 67.28 | 82.10 | 95.20 | 79.70 | 85.56 | 89.90 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
| Qwen2.5-1.5B | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
| Qwen2.5-3B | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
| Qwen2.5-7B | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
| Qwen3-4B | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
| Qwen3-30B | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
[…]
| GPT-5-mini | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
| Deepseek-V3.2 | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
| Gemini-2.5-Flash | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
[…]
| Qwen2.5-0.5B | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
| Qwen2.5-1.5B | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
| Qwen2.5-3B | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
| Qwen2.5-7B | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
[…]
| CRAFT | Trained | 4.75 | 10.09 | 7.95 | 20.20 | 30.14 | 20.85 | 20.50 | 25.65 | 12.50 |
| CRAFT | Trained | 18.30 | 26.89 | 22.57 | 44.18 | 57.84 | 55.25 | 42.86 | 48.34 | 42.29 |
| CRAFT | Trained | 34.11 | 43.71 | 61.20 | 57.24 | 71.15 | 92.60 | 54.90 | 63.47 | 80.10 |
| CRAFT | Trained | 54.65 | 64.59 | 81.40 | 66.09 | 81.04 | 95.70 | 76.88 | 83.34 | 93.20 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
| Qwen2.5-1.5B | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
| Qwen2.5-3B | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
| Qwen2.5-7B | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
| Qwen3-4B | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
| Qwen3-30B | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
[…]
| GPT-5-mini | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
| Deepseek-V3.2 | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
| Gemini-2.5-Flash | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
[…]
| Qwen2.5-0.5B | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
| Qwen2.5-1.5B | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
| Qwen2.5-3B | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
| Qwen2.5-7B | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
[…]
| CRAFT | Trained | 8.45 | 15.11 | 25.40 | 25.95 | 36.20 | 40.70 | 24.70 | 30.94 | 26.50 |
| CRAFT | Trained | 16.41 | 23.27 | 29.57 | 43.68 | 57.79 | 65.66 | 44.22 | 51.33 | 48.96 |
| CRAFT | Trained | 32.71 | 43.14 | 71.40 | 56.11 | 70.63 | 94.30 | 55.68 | 64.18 | 82.00 |
| CRAFT | Trained | 56.91 | 66.47 | 86.70 | 66.51 | 81.91 | 96.00 | 79.39 | 85.14 | 94.10 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | |||
| EM | F1 | EM | F1 | EM | F1 | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
| Qwen2.5-1.5B | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
| Qwen2.5-3B | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
| Qwen2.5-7B | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
| Qwen3-4B | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
| Qwen3-30B | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
[…]
| GPT-5-mini | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
| Deepseek-V3.2 | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
| Gemini-2.5-Flash | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
[…]
| Qwen2.5-0.5B | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
| Qwen2.5-1.5B | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
| Qwen2.5-3B | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
| Qwen2.5-7B | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
[…]
| CRAFT | Trained | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 | 0.05 |
| CRAFT | Trained | 22.95 | 31.91 | 45.70 | 59.04 | 38.40 | 46.19 |
| CRAFT | Trained | 31.80 | 42.07 | 53.15 | 67.10 | 49.30 | 57.19 |
| CRAFT | Trained | 56.25 | 65.38 | 63.25 | 78.00 | 67.15 | 73.77 |
| Model | Type | CRAFT | CRAFT | CRAFT | CRAFT | CRAFT | |||||
| Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | ||
[…]
| Avg. | 0.0 | 98.3(+98.3) | 1.1 | 99.2(+98.1) | 0.0 | 94.3(+94.3) | 1.0 | 99.4(+98.4) | 9.0 | 0.1(-8.9) | |
| HotpotQA | 0.0 | 99.6 | 0.1 | 99.8 | 0.1 | 97.5 | 0.8 | 99.8 | 5.7 | 0.1 | |
| 2WikiMQA | 0.0 | 99.5 | 3.2 | 99.7 | 0.0 | 92.5 | 2.3 | 99.8 | 16.5 | 0.1 | |
| MuSiQue | 0.0 | 95.8 | 0.0 | 98.2 | 0.0 | 92.8 | 0.0 | 98.5 | 4.9 | 0.0 | |
[…]
| Avg. | 0.0 | 99.3(+99.3) | 2.1 | 95.1(+93.0) | 0.2 | 99.5(+99.3) | 1.2 | 81.9(+80.7) | 0.8 | 100.0(+99.2) | |
| HotpotQA | 0.0 | 99.5 | 2.9 | 98.9 | 0.1 | 99.2 | 1.0 | 92.9 | 0.7 | 100.0 | |
| 2WikiMQA | 0.0 | 99.4 | 0.8 | 99.2 | 0.0 | 99.5 | 1.4 | 91.2 | 1.4 | 100.0 | |
| MuSiQue | 0.1 | 99.1 | 2.5 | 87.4 | 0.6 | 99.7 | 1.2 | 61.6 | 0.2 | 99.9 | |
[…]
| Avg. | 13.2 | 96.9(+83.7) | 50.3 | 98.8(+48.5) | 6.4 | 98.0(+91.6) | 15.6 | 98.1(+82.5) | 41.6 | 100.0(+58.4) | |
| HotpotQA | 14.8 | 98.2 | 55.2 | 99.2 | 4.3 | 99.4 | 27.6 | 99.2 | 32.8 | 100.0 | |
| 2WikiMQA | 7.7 | 99.0 | 33.8 | 99.4 | 6.3 | 99.1 | 8.2 | 98.8 | 23.9 | 100.0 | |
| MuSiQue | 17.2 | 93.5 | 61.8 | 97.6 | 8.7 | 95.6 | 10.8 | 96.3 | 68.1 | 100.0 | |
[…]
| Avg. | 94.1 | 99.3(+5.2) | 99.2 | 100.0(+0.8) | 94.4 | 97.2(+2.8) | 61.6 | 99.8(+38.2) | 15.4 | 100.0(+84.6) | |
| HotpotQA | 96.1 | 100.0 | 99.5 | 100.0 | 95.7 | 99.1 | 49.5 | 100.0 | 11.2 | 100.0 | |
| 2WikiMQA | 95.4 | 99.1 | 99.3 | 100.0 | 96.0 | 97.0 | 72.9 | 100.0 | 9.6 | 100.0 | |
| MuSiQue | 90.9 | 98.9 | 98.9 | 100.0 | 91.6 | 95.5 | 62.5 | 99.4 | 25.4 | 100.0 | |
Appendix B Additional Results
This section provides comprehensive supplementary results that expand on the main experiments in §4. We present training dynamics for all trace variants (§B.1), verification of faithfulness reward behavior (§B.2), full evaluation tables for CRAFT–CRAFT (§B.3), format compliance analysis (§B.4), format error patterns (§B.5), and model scale sensitivity (§B.6).
B.1 Training Dynamics for Additional Variants
Figures 5–8 show the training dynamics for CRAFT–CRAFT. Consistent with CRAFT (Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
Template-specific patterns: (1) CRAFT shows similar dynamics to CRAFT, with all four reward components contributing. (2) CRAFT/CRAFT lack but maintain stable training. (3) CRAFT uses only and ; the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
B.2 Faithfulness Reward Verification
As noted in §4.4, the faithfulness reward remains low during training, particularly for smaller models. To verify that this reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0). Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%. This indicates that while 69.3% of correct predictions exhibit fully faithful reasoning, 30.7% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness. This test-time value is approximately six times higher than the training-time average of 0.11, confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
B.3 Full Results for Additional Variants
Tables A–A present the complete evaluation results for CRAFT–CRAFT across all model scales and baselines, complementing the CRAFT results in the main paper (Table 3.3.2).
CRAFT Results.
CRAFT retains the same structure as CRAFT but removes <plan>. CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base). Notably, CRAFT surpasses all API models on MuSiQue and 2WikiMHQA while achieving competitive Faithfulness scores (78.35%–95.20%).
CRAFT Results.
CRAFT removes explicit document citation (<gold_docs>), testing whether models can maintain faithfulness without explicit grounding signals. CRAFT achieves 54.65% EM on MuSiQue, 66.09% on HotpotQA, and 76.88% on 2WikiMHQA. Despite lacking supervision, the model maintains high Faithfulness (81.40%–95.70%), suggesting that plan-reason-answer consistency provides sufficient structure for faithful reasoning.
CRAFT Results.
CRAFT retains only <reason>+<answer>. CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA. EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance with appropriate reward design.
CRAFT Results.
CRAFT uses direct answer generation without structured reasoning traces ( and not applicable). CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA. A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training.
B.4 Format Compliance Analysis
Expanding on the format compliance summary in §4.4, Table A presents a comprehensive analysis across all CRAFT variants, model scales, and individual datasets.
GRPO substantially improves format compliance across all scales. Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates, while even 3B models only reach 6.4%–50.3% on average. GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases. Even 7B base models benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%). This demonstrates that successfully teaches models to follow structured output requirements across all capacity levels.
Smaller models struggle with complex templates but benefit most from GRPO training. For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a notable failure on CRAFT, where GRPO format compliance drops from 9.0% to 0.1%. We attribute this to a reward signal collapse: CRAFT uses only and , but the 0.5B model lacks sufficient capacity to produce correct answers without reasoning structure, yielding near-zero throughout training. With as the only effective reward, the model receives no learning signal to distinguish between the minimal <answer> format and the more complex XML structures encountered during exploration, leading to format drift rather than convergence. This failure is consistent with the complete accuracy collapse of CRAFT on CRAFT (Table A). In contrast, larger models master all formats: 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
| Version | Base Model Errors (7B) | GRPO Model Errors (7B) |
[…]
Missing closing tags (60%):
<answer>1958 Award missing </answer> Truncation (30%): response cut mid-sentence |
Empty answer (rare):
<answer> no content |
[…]
Missing closing tags (70%):
<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): </answer> used for </reason> |
No errors found ✓ |
[…]
Missing answer block (45%):
entire <answer>...</answer> absent Truncation (40%): response cut off |
Malformed tags (3%):
truncated closing tag |
[…]
Missing closing tags (85%):
<answer>Russell Humphreys missing </answer> |
Extra text (0.2%):
parenthetical note after tags |
[…]
Wrong format (85%):
bare text without XML tags e.g., A Bug’s Life |
No errors found ✓ |
B.5 Format Error Analysis
Table 11 complements the quantitative results in Table A by presenting representative error patterns across all CRAFT variants. Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
B.6 CRAFT Template Sensitivity
As observed in Table 3.3.2, CRAFT shows limited but non-trivial gains at the smallest scale. Here we examine this model’s sensitivity to trace template choice in detail, revealing striking differences between SFT and GRPO failure modes. SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%). In contrast, GRPO rescues templates CRAFT–CRAFT but still fails on CRAFT. Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
This dichotomy reveals three key insights: (1) SFT cold-start is fragile for 0.5B models: templates with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse. (2) GRPO’s exploration mechanism overcomes SFT’s brittleness: reinforcement learning enables the model to discover valid reasoning paths even on complex structures, as long as some structural scaffolding exists. (3) CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure for stable learning. The optimal template for 0.5B deployment is CRAFT (<reason>+<answer>), providing the best balance of structure and simplicity.
| Model | Judge | MuSiQue | HotpotQA | 2WikiMHQA | |||
| Faith. | Faith. | Faith. | |||||
| Qwen2.5-7B | Qwen3-Max | 58.25 | – | 83.65 | – | 76.85 | – |
| Gemini-Flash | 55.40 | -2.85 | 82.15 | -1.50 | 73.90 | -2.95 | |
| Qwen2.5-7B | Qwen3-Max | 50.85 | – | 83.50 | – | 78.95 | – |
| Gemini-Flash | 49.70 | -1.15 | 81.40 | -2.10 | 77.50 | -1.45 | |
| CRAFT w/o judge | Qwen3-Max | 71.20 | – | 91.80 | – | 89.25 | – |
| Gemini-Flash | 69.85 | -1.35 | 90.25 | -1.55 | 86.80 | -2.45 | |
| CRAFT (full) | Qwen3-Max | 80.35 | – | 95.00 | – | 94.80 | – |
| Gemini-Flash | 77.60 | -2.75 | 93.60 | -1.40 | 93.25 | -1.55 | |
| Metric | MuSiQue | HotpotQA | 2WikiMHQA |
| Pearson | 0.998 | 0.999 | 0.997 |
| Spearman | 1.000 | 1.000 | 1.000 |
| Mean | 2.03 | 1.64 | 2.10 |
| Ranking Agreement | 4/4 | 4/4 | 4/4 |
Appendix C Cross-Judge Evaluation
To address the potential circularity of using Qwen3-Max as both the training reward provider and evaluation judge, we re-evaluate all model configurations using Gemini-2.5-Flash Google DeepMind (2025) as an independent judge with identical evaluation prompts. Gemini-2.5-Flash is not involved in any stage of CRAFT training.
Table 12 reports the faithfulness scores from both judges across all model configurations under CRAFT. Gemini-2.5-Flash assigns systematically lower absolute scores than Qwen3-Max (mean = 1.92), which is expected given known differences in evaluation stringency across models. Crucially, the relative ordering of all four configurations is fully consistent under both judges across all three benchmarks (Spearman = 1.0).
Table 13 confirms strong agreement: Pearson and Spearman across all benchmarks, demonstrating that the reported faithfulness improvements are robust to judge selection.
Appendix D Case Studies
| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
[…]
Subq1: Where is the Oh Yeah album from?
Subq2: What is the largest city and capital of that state? Subq3: Who won the Indy car race in that city? |
Subq1: Where is the Oh Yeah performer from?
Subq2: What is the largest city of that state? Subq3: Who won the Indy car race there? |
[…]
| 1 - Plan steps followed in order | 1 - Plan followed | |
| 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs in gold_docs | |
| 1 - Answer matches reasoning conclusion | 1 - Mario Andretti supported | |
| 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified | |
| 0.50 | 1.0 | |
We present an additional case from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs. This case complements the right-answer-wrong-reason example in §4.6 by showing a case where both accuracy and faithfulness are corrected.
D.1 Case: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
Question: Who won the Indy car race in the largest city of the state where the Oh Yeah album performer was born?
Gold Answer: Mario Andretti